Grade 12Physical Chemistry

For a solution of 0.849 g of mercurous chloride in 50 g of HgCl2. [L] the freezing point depression is 1.24ºC. Kf for HgCl2 is 34.3. What is the state of mercurous chloride in HgCl2 ? (Hg – 200, Cl – 35.5)

AKSHAY CHAUHAN
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
delta T=change in freezing point of the solution= Kf * molality.............(1)
now calculate the molality of the solute(Hg2Cl2).
To calculate the molality, calculate the number of mole of mercurus chloride.
moles=0.849/M
where M=molecular weight ofmercurus chloride.
this amount is in 50 g of HgCl2
so, in 1000g, moles of solute=0.849/M*1000=molality
Now apply (1) and calculate the decrease n freezing point of the solution.
Now if we have the information of normal freezing point of HgCl2, we can also get the state of solution.